OpenStudy (anonymous):

Solve each equation using logarithms. Round to the nearest ten-thousandth. 1) 3^b= 17 2) 12^r= 13 3) 5x18^6x= 26 4) 11^n-8-5= 54 5) 16^n-7+5= 24

OpenStudy (anonymous):

I'll show the first one. You should be able to do the second one after that. \[3^b = 17\] Take the log of both sides (You can choose base 3, but you don't have to. I'm going to choose base 10) \[Log(3^b) = Log(17)\] Now, use your log rules and solve for b: \[b Log(3) = Log(10)\] \[b = \frac{Log(10)}{Log(3)}\] Just throw that into a calculator :) Can you do 2?

OpenStudy (anonymous):

Oh, for some reason I changed my 17 to a 10. Sorry about that. \[b = \frac{Log(17)}{Log(3)}\]